### v1

# Predictive Modeling of Power Outages during Extreme Events: Integrating Weather and Socio-Economic Factors

Abstract
This paper presents a novel learning-based framework for predicting power outages caused by extreme events. The proposed approach specifically targets low-probability, high-consequence outage scenarios and leverages a comprehensive set of features derived from publicly available data sources. We integrate EAGLE-I outage records (2014–2024) with weather, socio-economic, infrastructure, and seasonal event data. Incorporating social and demographic indicators reveals underlying patterns of community vulnerability and provides a clearer understanding of outage risk during extreme conditions. Four machine learning models—Random Forest (RF), Support Vector Machine (SVM), Adaptive Boosting (AdaBoost), and Long Short-Term Memory (LSTM) —are evaluated. Experimental validation is performed on a large-scale dataset covering counties in the lower peninsula of Michigan. Among all models tested, the LSTM network achieves the lowest prediction error. Additionally, the result demonstrates that stronger economic conditions and more developed infrastructure are associated with lower outage occurrence.
keywords:
Power system resilience, power outage prediction, machine learning.
1 Introduction
Low-Probability High-Consequence (LPHC) events—such as floods, hurricanes, and winter storms—pose severe threats to power systems reliability and resilience. [1]. These events can cause large-scale blackouts and economic losses, as seen in Hurricane Sandy (2012), Irma (2017), and the 2021 Texas winter storm [2, 3, 4]. Several studies has explored a wide range of modeling approaches to understand and predict the relationship between extreme weather conditions and outage occurrence.
For example, Eskandarpour et al. [5] predicted component damage prior to hurricanes by establishing decision boundaries using logistic regression. Spatial–temporal poisson regression has also been applied to model transmission line failure probabilities during severe weather events [6]. Although these methods offer clear interpretability, their linear assumptions limit their ability to capture the complex multi-variable interactions.
ML models provide an effective framework to manage complex non-linear relationships and process substantial data volumes in real-time [7]. These models outperform traditional statistical approaches due to their ability to model nonlinearities and integrate diverse datasets, including weather, infrastructure, and socio-economic indicators. Classification-based ML studies such as [8, 9] used models like SVM and boosted classifiers to determine outage occurrence or component damage. Regression-based studies have used tree-based ensembles such as Decision Trees, RF, AdaBoost, and XGBoost to predict outage counts, durations, or affected customers [10, 11, 12]. Deep learning approaches further extend predictive capability by capturing temporal and sequential patterns in outage and weather data. Arif et al. [13] employed deep neural networks to estimate restoration times, and recurrent architectures such as recurrent neural networks (RNN) to model temporal outage progression under varying environmental conditions. These models offer strong predictive performance but require large, high-quality datasets. Some studies examined on the outages driven by routine weather or equipment failures. Jaech et al. [14] applied RNN-based models with environmental and operational data to predict outage duration in Seattle, while Sun et al. [15] used AdaBoost to estimate outage counts in Kansas. Sarwat et al. [16] showed that day-to-day weather conditions influence power outage. Our work focuses on LPHC events that generate large-scale disruptions and require fundamentally different modeling considerations.
The state-of-the-art studies often encounter challenges that can compromise the accuracy of their analyses. A major limitation is the scarcity of historical data on extreme events and the associated power outages. In particular, during such events, several monitoring devices may fail, resulting in incomplete and imbalanced datasets. In [17, 18, 19], Synthetic Minority Oversampling Technique (SMOTE) has been used to address data imbalance in classification tasks in power outage studies. However, SMOTE is effective for categorical targets, it is less productive for LPHC outage prediction, which is framed as a regression problem. Another limitations arises when models are trained on a single extreme event, as they exhibit limited adaptability to concurrent incidents with complex, multi-variable interactions [20]. As shown in [21], combining data from two winter storm events instead of one improves power outage prediction. Furthermore, socio-economic features are often missing in the previous models. As demonstrated in Cruz et al. [11], socio-economic indicators such as income levels and housing characteristics provide important context for understanding community vulnerability. Counties with lower income or higher unemployment tend to have fewer resources for preparedness, which can increase outage severity and prolong restoration times. Similar observations have been reported in other studies [13, 14], where result shows that socio-economic factors contribute to prolonged outage durations and higher customer outage number. Incorporating socio-economic variables therefore enables a more comprehensive understanding of outage dynamics, supports improved prediction performance for LPHC events.
In our prior work [19], we developed ML models to predict power outage probabilities, investigating the impact of socio-economic and infrastructure factors on non-LPHC outage events. This paper builds on those findings and proposes enhanced methodologies to predict power outages due to extreme weather events. The contributions of this paper are outlined as follows:
1) Comprehensive Data Integration: Our model combines weather, socio-economic, and infrastructure data, along with seasonal indicators, to capture the complex factors influencing power outages. This comprehensive integration significantly enhances the predictive performance of the model.
2) Innovative Use of SMOGN and KNN to Address Data Scarcity in LPHC Events: This work introduces the application of the Synthetic Minority Over-sampling Technique for Regression (SMOGN) and K-Nearest Neighbors (KNN) methods to mitigate data imbalance and missing data challenges associated with LPHC events. By adapting these techniques to the outage prediction context, the proposed framework effectively addresses data scarcity and improves the accuracy and robustness of LPHC event prediction.
3) Evaluation of ML Models: We evaluate multiple ML and deep learning (DL) models, with LSTM proven particularly effective due to its ability to capture temporal patterns in time-series data. This makes our model highly suited for real-world power outage forecasting.
4) Large-Scale Validation: We validate our approach on a large-scale dataset spanning 2014 to 2022 across Michigan counties, ensuring the model’s robustness and generalizability to diverse geographic regions and socio-economic conditions.
The remainder of the paper is structured as follows: Section II presents the data description and preparation. In section III, the methodology is discussed. Experimental setup is presented in section IV. The experimental results and discussion are provided in section V. Finally, the conclusion is drawn in section VI.
2 Data Description and Preparation
2.1 Historical Outage Data
In this study, the EAGLE-I dataset is utilized to analyze severe power outages caused by LPHC events. The EAGLE-I dataset [22], collected and managed by Oak Ridge National Laboratory (ORNL), spans from November 2014 to December 2024. This dataset, updated every 15 minutes, aggregates data from utility companies’ public outage maps using an ETL (Extract, Transform, Load) process to estimate the number of customers affected by power outages in each county. For this analysis, we focus on the outage data relevant to Michigan. Fig. 1 illustrates the daily number of customer outages in Wayne County throughout 2024.
Figure 1: Number of customers outage in Wayne County during 2024 based on EAGLE-1 dataset.
In this context, a “customer” is defined as any entity that purchases energy from a utility under a specific tariff, rather than individual persons. Given that in the EAGLE-I, number of customers affected by power outage is recorded, and to ensure a fair comparison between counties with different sizes, it is crucial to adjust it for population differences. This can be done by collecting population data for each county from the U.S. Census Bureau [^2] https://data.census.gov/table. This approach allows us to normalize outage data relative to population size. In this way, the number of customers in the EAGLE-I dataset is divided by the population of the county scaled by 2 as shown in (1).
$P_{o}=\frac{Number\>of\>customers\>lost\>power\>*2}{Population\>of\>the\>county}$ (1)
It is estimated that, on average, each customer corresponds to approximately two residents [23].
2.2 Historical Weather Data
Historical hourly weather data was obtained from Open-Meteo, an open-source weather API that provides high-resolution meteorological information on an hourly basis  [^3] https:https://open-meteo.com/. This data source serves as the dynamic variable of the model. For instance, weather data offers insights into conditions that could lead to outages, such as wind speed and temperature. We have chosen 8 features including air temperature (F), precipitation (inch), wind speed (km/h), wind gusts (km/h), shortwave radiation ($W/m^{2}$), relative humidity (%), cloud cover (%), and surface pressure (hPa). The Fig. 2 shows the locations of weather stations across the Lower Peninsula of Michigan used in this study.
The historical weather data often has missing data, particularly during LPHC events. To address this issue, we employ K-Nearest Neighbors (KNN) imputation method describe as follows:
Algorithm 1 KNN Algorithm for Finding the Nearest Counties
Input:
$k\leftarrow 5$ (Number of nearest counties to be found)
$S_{i}\leftarrow\text{Station}_{i}$ (Find the k nearest counties of $S_{i}$)
$(x_{i},y_{i})$ (Coordinates of $S_{i}$)
$D_{ij}$ (Euclidean distance between $S_{i}$ and $S_{j}$)
Procedure:
for $k=1,\dots,j$ do
      Calculate $D_{ij}$ between $S_{i}$ and $S_{j}$ using equation (2)
      Sort distances between $S_{i}$ and all other counties in descending order
      Select the first $k$ closest counties to $S_{i}$
end for
Output: 5 nearest counties to $S_{i}$
For a given $county_{i}$, the $k$ closest counties are determined by calculating the Euclidean distance between $county_{i}$ and every other county, as described in (2).
$\text{Euclidean Distance}(S_{i},S_{j})$ $=\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}}$ (2)
where $(x_{i},y_{i})$ and $(x_{j},y_{j})$ represent the coordinates of counties $S_{i}$ and $S_{j}$, respectively. Then, impute the missing data by averaging the selected neighbors:
$Missing\>Value=\frac{1}{j}\sum_{i=1}^{j}{Observed\>value_{i}}$ (3)
where $j$ represents the number of nearest counties considered for the imputation, identified by the K-Nearest Neighbors (KNN) algorithm [24]. Finally, the dataset is normalized using min-max scaler.
Figure 2: Weather stations locations in Lower Peninsula of Michigan.
2.3 Socio-economic Data
Socio-economic indicators provide meaningful insights into community vulnerability and resilience. This hidden knowledge is instrumental in understanding factors that are not immediately obvious from basic data analysis. For this study, socio-economic indicators were obtained from the U.S. Census Bureau, including the distribution of the year when residential structures were built, the average household income, and the county-level unemployment rate. The selected data was retrieved from the U.S. Census Bureau American Community Survey (ACS) 5-Year Estimates via the Census API.  [^4] https://api.census.gov/data/2021/acs/acs5
The distribution of the year when residential structures were built reflects the age and durability of the electric infrastructure. Average household income and unemployment capture the economic capacity and stress of a community, influencing both infrastructure investment and the speed of recovery from extreme weather events. Together, these factors complement power infrastructure data by capturing socio-economic conditions that influence outage occurrence and restoration.
2.4 Power Infrastructure Data
The infrastructure features, including the number of poles, towers, substations, transformers, and lines for each county, are collected from the Open Street Map dataset [25] as shown in Fig. 3. Given that the detailed topology of the power distribution network is typically not publicly available, data from the Open Street Map can serve as a useful approximation of the actual topology for research purposes. Furthermore, they provide information on how extensively the system might be affected by extreme conditions such as high winds [26].
For this study, the total number of each feature within a county is aggregated and then counted to provide a comprehensive overview of the area’s characteristics. Then, we perform column-wise normalization, where we divide the total number of each type of infrastructure in a county by the sum of that feature across the entire state. This method allows us to determine which county has a higher proportion of a particular infrastructure component, such as towers.
Figure 3: Power infrastructure network of counties in Lower Peninsula of Michigan.
3 Data-driven Outage Prediction Algorithm
The proposed multi-step county-level power outage prediction algorithm is illustrated in Fig. 4. The algorithm follows a four-step process. First, data preparation is performed, where different input features are preprocessed and integrated into a unified dataset. Second, the integrated dataset is rebalanced to reduce the impact of data imbalance and missing data. Third, multiple machine learning models are trained to capture the complex relationships between weather, socio-economic, and infrastructure factors and outage occurrence. Finally, the performance of the models is evaluated using standard metrics.
Figure 4: The framework of the proposed power outage prediction algorithm.
3.1 Identifying Extreme Weather Event Data
The outage and weather dataset spans the period from 1 November 2014 to 30 December 2024. Within this time frame, a total of 236 weather-related extreme outage events were identified from the National Centers for Environmental Information (NOAA)  [^5] https://www.ncdc.noaa.gov/stormevents/. Fig. 5 illustrates the distribution of extreme weather events across 20 Michigan counties.
Figure 5: Number of Extreme Events by County (2014–2024).
This distribution shows that extreme events are sparse. To address this limitation, seasonal information aligned with the timing of distinct extreme weather patterns is incorporated. For a specified extreme event day (denoted as $d_{ex}$), we identify all days within the same season and include them in the set $R$. This is defined as:
$R=\{d_{1},d_{2},\ldots,d_{j}\}\text{ where }d_{j}\in(m_{i-1},m_{i},m_{i+1})$
where, $m_{i}$ represents a month within a season, selected from the set $M=\{m_{1},m_{2},m_{3},\ldots,m_{12}\}$.
The next step involves calculating the $L2norm$ of the weather features between the extreme day ($d_{ex}$) and all other days in set $R$. This comparison helps identify days with weather patterns most similar to the extreme event day.
Assume $w_{d_{ex}}$ is the matrix of weather features for the extreme day (comprising 8 features, one for each weather features). The weather features for the $j$ days are represented by the matrix $W_{d}$ as follows:
$W_{d_{ex}}=\begin{pmatrix}w_{1d_{ex}}&w_{2d_{ex}}&\cdots&w_{8d_{ex}}\\ \end{pmatrix}$
$W_{d}=\begin{pmatrix}w_{1d_{1}}&w_{2d_{1}}&\cdots&w_{8d_{1}}\\ w_{1d_{2}}&w_{2d_{2}}&\cdots&w_{8d_{2}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{1d_{j}}&w_{2d_{j}}&\cdots&w_{8d_{j}}\\ \end{pmatrix}$ (4)
The L2normL2\>norm between the row vector $w_{d_{ex}}$ of $W_{d_{ex}}$ and each row vector $w_{d_{j}}$ of $W_{d}$ is calculated using the following formula:
$L2\>norm(w_{d_{ex}},w_{d_{j}})=\sqrt{\sum_{k=1}^{8}(w_{kd_{ex}}-w_{{k}d_{j}})^{2}}$ (5)
Here, $W_{d_{ex}}=\begin{pmatrix}w_{1d_{ex}}&w_{2d_{ex}}&\cdots&w_{8d_{ex}}\\ \end{pmatrix}$ is the row vector of extreme day’s weather features, $W_{d_{j}}=\begin{pmatrix}w_{1d_{j}}&w_{2d_{j}}&\cdots&w_{8d_{j}}\\ \end{pmatrix}$ is the j-th row vector of $W_{d}$, $w_{kd_{ex}}$ and $w_{{k}d_{j}}$ are the elements of $W_{d_{ex}}$ and $W_{d_{j}}$, respectively.
Subsequently, the calculation is repeated for each row in the matrix $W_{d}$, corresponding to each day $d_{j}$ within the set $R$. For each day, compute the L2normL2\>norm between the weather feature vector of the extreme day $W_{d_{ex}}$ and that of the day $d_{j}$, denoted as $W_{d_{j}}$. This results in a $j\times 1$ matrix where each element represents the $L2norm$ between $W_{d_{ex}}$ and $W_{d_{j}}$. The process is outlined mathematically as follows:
$L2\text{Norm}=\begin{pmatrix}L2\>norm(W_{d_{ex}},W_{d_{1}})\\ L2\>norm(W_{d_{ex}},W_{d_{2}})\\ \vdots\\ L2\>norm(W_{d_{ex}},W_{d_{j}})\\ \end{pmatrix}$ (6)
Finally, we select the weather features and outage data of days with the lowest $L2norm$ from matrix $L2Norm$, indicating the highest similarity to the extreme event day. Incorporating these seasonal references significantly enhances the model’s ability to differentiate between various outages in the dataset. Then, the dataset is rebalanced and used to train the machine learning models described in the following subsections.
Algorithm 2 SMOGN for Outage-Ratio Rebalancing
Input:
    $D\leftarrow$ Training dataset (hourly), with features X
    $T\leftarrow$ Threshold of minority and majority instances
    $k\leftarrow$ Number of nearest neighbors
    $O\%\leftarrow$ Percentage of over-sampling
    $U\%\leftarrow$ Percentage of under-sampling
Partition:
    $S_{min}\leftarrow$ $y_{i}\geq T$ (Minority instances)
    $S_{maj}\leftarrow$ $y_{i}<T$ (Majority instances)
/*Over-Sampling minority instances*/
for $S\in S_{min}$ do
  $\hat{S}_{min}\leftarrow|S|\times O\%$
  for $sample\in S$ do
   $KList\leftarrow$ $k$ nearest neighbors of sample
   $Distances\leftarrow$ distance between sample and each neighbor
   $SafeRange\leftarrow{0.5\times\text{median}(Distances)}$
   for $i=1$ to $\hat{S}_{min}$ do
     $X\leftarrow$ randomly selected neighbor from KListKList
     if $Distance(X,sample)<SafeRange$ then
      $X_{new}\leftarrow$ SMOTER interpolation
     else
      $X_{new}\leftarrow$ samplesample + Gaussian Noise
     end if
     Append $X_{new}$ to $D_{new}$
   end for
  end for
end for
/*Under-Sampling majority cases*/
for $S\in S_{maj}$ do
  $under\_sampled\_instances\leftarrow$ randomly sample $|S|\times U\%$ instances
  Append under_sampled_instancesunder\_sampled\_instances to $D_{new}$
end for
Output: Rebalanced training set $D_{new}$
3.2 Input Data Balancing
A key challenge in predicting LPHC outages is the imbalance in outage data, where small-scale events dominate and large-scale extreme events are minority. This imbalance causes ML models to favor frequent low-impact cases, leading to poor performance on the critical but infrequent LPHC events. In order to balance the dataset, we apply SMOGN, which has been introduced to address the issue of imbalanced datasets in regression tasks [27]. It integrates two over-sampling technique, SMOTER and SMOTER-GN [28]. While the $L2norm$ filtering step extracts outages associated with extreme events, the SMOGN algorithm addresses the imbalance in outage severity within these events. SMOGN generates synthetic data by applying three key techniques: random under-sampling, SMOTER (Synthetic Minority Over-sampling Technique for Regression), and adding Gaussian noise. Random under-sampling involves randomly removing samples from the majority class. It generates new synthetic samples by interpolating rare cases. This is done by taking a weighted average of the target variable values between the original minority instance and its k-nearest neighbors [28].
Figure 6: Probability distribution of customer outages in the EAGLE-I dataset (2014–2024), shown before and after applying SMOGN.
SMOGN is summarized in Algorithm 2: We mark a case as rare if its outage value is in the top 30% of the target distribution. A rarity threshold of 380 customers out was used and five nearest rare dataset neighbors were considered for SMOTER interpolation within a safe region. Also, Gaussian noise with perturbation 2% was added when neighbors lie outside the local safe range. Minority cases were oversampled approximately one-to-one, and the majority was undersampled to one half. Fig. 6 shows the influence of SMOGN on the outage occurrence probability distribution for the extreme weather dataset used in the case study.
The original dataset is dominated by small outages near zero, while the rebalanced dataset includes more high-consequence outages. After rebalancing, probability is reduced in the low-outage region and increased in the mid/high-outage region beyond the threshold.
3.3 Model Selection and Training
After the input data has been rebalanced, the following ML models are trained to capture its complex patterns:
Random Forest (RF): RF for regression is an ensemble learning method that constructs multiple decision trees during training and outputs the average prediction of all trees for regression tasks. Each tree is trained on randomly selected subsets of data, reducing the correlation between trees and mitigating over-fitting. To predict customer outages number, a Random Forest Regressor was trained using weather and socio-economic features. The model is trained with 100 decision trees and evaluated on the same dataset.
Support Vector Machine (SVM): SVM for regression identifies an optimal hyperplane in a higher-dimensional space that best fits the data while maximizing the distance from the nearest points. It minimizes errors, making it suitable for predicting continuous variables. SVM uses kernel functions to map the data into a space where these nonlinear patterns become easier to fit.
Adaptive Boosting (AdaBoost): It is a boosting algorithm for regression which strengthens weak models by prioritizing harder-to-predict instances, refining the ensemble’s accuracy with each iteration. The final prediction is a weighted sum of the predictions from all weak learners, with weights determined based on their performance. Given the imbalanced nature of outage data, AdaBoost can concentrate learning on rare and high-impact outage conditions.
Long Short-Term Memory (LSTM): LSTM networks are a type of RNN specialized in processing sequences of data. They are particularly effective for regression tasks involving time-series data due to their ability to capture temporal dependencies over different time intervals. In this study, historical weather, socio-economic, and infrastructure features are processed through the LSTM to capture patterns over time, followed by a dense output layer for outage prediction.
3.4 Performance Evaluation
The following common performance metrics are used for models evaluation [29]:
Mean Absolute Error (MAE): It measures the average deviation of the predicted values from the actual observed values. It shows how far the model’s predictions are from the actual values on average.
[…]
Mean Square Error (MSE): That metric measures the average of the squared differences between the estimated values and the actual observed values. Lower MSE indicates that the model is more consistent and accurate across the full range of observations.
$MSE=\frac{1}{n}\sum_{i=1}^{n}(P_{o}(i)-\hat{P_{o}(i)})^{2}$ (8)
In both expressions, $P_{o}(i)$ denotes the actual observed value and $\hat{P_{o}(i)}$ represents the predicted value.
3.5 Outage Prediction Model Formulation
With all relevant features integrated into the dataset, the problem is formulated as predicting county-level power outages under extreme events. Eq. 3.5 shows the model formulation.
$P_{o}(t)$ $=f\big(P_{o}(t-1),P_{o}(t-2),\cdots,P_{o}(t-n),$
$\quad W(t),W(t-1),\cdots,W(t-n),S,I_{f}\big)$ (9)
Here, $P_{o}(t)$ denotes the ratio of customers without power in a county at time $t$ during an extreme event, and $P_{o}(t-k)$ is the corresponding ratio observed during the $k$-th most recent extreme event. $W(t)$ is the vector of contemporaneous weather features for the county, while W(tk)W(t-k) contains the weather features recorded for the $k$-th most recent event. $S$ comprises socio-economic indicators, including average household income, unemployment rate, and the distribution of year built for residential structures. $I_{f}$ aggregates power-system attributes such as counts of poles, towers, substations, transformers, and line segments by county.
4 Experimental Setup and Simulation Results
The dataset is divided into training (2014 to 2018), validation (2019), and test (2020–2021) sets to evaluate the performance of the models. The implementation of RF, SVM, and AdaBoost is done using Python Scikit-learn library [30] and LSTM is implemented using Tensorflow library [31].
We conducted hyperparameter tuning utilizing the Grid- Search tool in Python [30], employing cross-validation to identify the optimal input parameters for the models, as shown in Table 1. The LSTM network is a one-layer LSTM with 128 nodes followed by a Dense layer. The LSTM layer employs its standard activation structure, where the cell state is updated using a tanh activation and the gating mechanisms use sigmoid functions. Adam optimizer is used with a learning rate of 0.001 and a batch size of 128.
Table 1: Cross-Validation Hyper-Parameters Selection
Hyper-Parameter RF SVM AdaBoost
Kernel rbf
Learning Rate - - 0.001
Estimators 4 - 120
Loss square
C - 10 -
Max Depth 100 - -
Gamma - 0.001 -
Fig. 7 illustrates the correlation analysis between socio-economic, infrastructure features and power outages. Higher mean income, lower unemployment rate, and more extensive infrastructure, including towers, lines, transformers, and poles, are associated with lower outages, suggesting that wealthier counties or those with more infrastructure might have higher power systems resilience. Substations and the year of structure construction also show a slight correlation, indicating newer or more substations may lead to improved outage resistance.
Figure 7: Pearson’s correlation of the customers outage number with the infrastructure and socio-economic variables.
We also conducted a feature importance analysis using the RF-based Mean Decrease in Impurity (MDI). Fig. 8 depicts the contribution of each feature to the model’s predictive accuracy. The analysis showed that weather variables dominate outage prediction, with precipitation (19%), wind speed (13.5%), and surface pressure (13.1%) being the top contributors. Non-weather features, including infrastructure density (3.5%) and socioeconomic indicators such as average income (1.4%), also exhibit contribution to the prediction.
Figure 8: Features importance across weather, infrastructure, and socioeconomic variables.
After these relationships are established, the models are trained using different combinations of input data, both with and without the inclusion of socioeconomic, infrastructure, and weather features, so that the influence of each data layer on predictive performance can be examined. By comparing model accuracy across these configurations, the added value of these additional features is evaluated, and their overall impact on the outage prediction process is determined.
Scenario-1 $=\big(P_{o}(t-1),P_{o}(t-2),\cdots,P_{o}(t-n),$
$\quad W(t),W(t-1),\cdots,W(t-n)\big)$
Scenario-2 $=\big(P_{o}(t-1),P_{o}(t-2),\cdots,P_{o}(t-n),$
$\quad W(t),W(t-1),\cdots,W(t-n),S,I_{f}\big)$
Scenario 1 uses weather data from the event day and the preceding day, together with outage information from the day before the extreme event and seasonal weather patterns. This configuration enables the models to utilize historical outage behavior along with weather conditions to estimate future outages. Scenario 2 expands upon this by incorporating all data used in Scenario 1 together with infrastructure and socioeconomic information.
Table 2: Performance Comparison of Models in Scenario-1 and Scenario-2.
Model Scenario-1 Scenario-2
MSE MAE MSE MAE
RF 0.0231 0.0898 0.0222 0.0827
SVM 0.0175 0.0796 0.0172 0.0774
AdaBoost 0.0210 0.0735 0.0204 0.0730
LSTM 0.0127 0.0749 0.0086 0.0432
Figure 9: Performance comparison between models in terms of MSE.
Table. 2 presents the performance of each model and shows that the scenario incorporating socioeconomic and infrastructure information produces lower prediction error. This outcome provides evidence that incorporating socioeconomic and infrastructure variables enhances the model’s predictive capability.
Fig. 10 illustrates the performance of the model by comparing the predicted outage values with the actual recorded values for LSTM model. Each point represents a data sample from the validation period, and the diagonal line indicates the ideal one to one correspondence between predicted and observed outages. As shown in the figure, the majority of the predictions follow the general upward trend of the reference line.
Figure 10: Scatter-plots of predicted versus actual customer outage for LSTM model.
Fig. 11 illustrates the comparison between the predicted and actual customer outage counts for the June 6–8, 2020 flood event in Wayne County, Michigan using four models: RF, SVM, AdaBoost, and LSTM. The time series plots show how well each model follows the temporal evolution and magnitude of the observed outages. Among the four models, the LSTM provides the closest match to the actual outage pattern. This indicates that the LSTM is able to effectively exploit temporal dependencies in the sequence of weather and related features during the event window. The Random Forest model also exhibits good performance. The AdaBoost model shows moderate agreement with the actual outages. The SVM model presents the weakest performance among the four.
The performance of our proposed outage prediction framework is also compared with state-of-the-art studies in Table 3. To ensure fairness, the comparison is limited to works that employed the EAGLE-I dataset as the primary source of historical outage records. Table 3 highlights that prior studies either rely on limited feature sets or provide low-resolution prediction outputs, such as maximum or averages outage customers number. In contrast, our method integrates weather, socio-economic, infrastructure, and seasonal indicators to deliver accurate county-level, hourly outage predictions during extreme events.
(a) RF prediction vs. actual outages
(b) SVM prediction vs. actual outages
(c) AdaBoost prediction vs. actual outages
(d) LSTM prediction vs. actual outages
Figure 11: Comparison of predicted and actual outage counts for the June 2020 flood event in Wayne County.
Table 3: Comparison of Proposed Work With State-of-the-Art Studies.
Study Dataset Features Used Best Model Prediction Resolution
Lee et al. [12] EAGLE-I + NWS alerts Weather features only XGBoost
($R^{2}$ = 0.998, RMSE = 2115 customers)
Event-level maximum outage prediction;
No hourly or county-level resolution
Cruz et al. [11] EAGLE-I + NASA weather + Census data Weather + Socio-economic RF: MSE = 5.93e-05
MAE = 0.00102
RMSE = 0.00770
Next-day average outage prediction;
Not high-resolution
Our Work (Proposed) EAGLE-I + Open-Meteo + Census + Infrastructure Weather + Socio-economic + Infrastructure + Seasonal indicators LSTM
(MSE = 0.0086, MAE = 0.0432)
County-level, hourly outage prediction during extreme events
Given that this model is primarily trained on data from extreme weather and socio-economic features, it is well-suited to be used for planning to enhance the resilience of power systems. Planning can be done through simulating the system’s performance across a range of severe weather scenarios.
5 Conclusions
In this paper, we presented a novel machine learning–based approach for predicting power outages caused by LPHC events. The method integrated historical outage data from the EAGLE-I dataset, weather data from Open-Meteo, as well as infrastructure and socio-economic information to construct predictive models. We addressed the challenges of limited data availability and missing records for LPHC events by applying KNN and SMOGN methods. We showed the correlations between socio-economic and infrastructure features with extreme power outages. Several machine learning models, including RF, SVM, AdaBoost, and LSTM networks, were compared. The experimental evaluation on Michigan counties demonstrated that the LSTM model consistently outperformed RF, AdaBoost, and SVM in terms of lowest error. Our analysis indicated that weather-related variables—particularly precipitation and wind speed were the primary drivers of outages, while infrastructure and socio-economic factors exerted secondary influences. This model can be adapted to other regions by incorporating localized data, including region-specific weather patterns and infrastructure conditions as well as local social-economic information. Similarly, extending the model to handle other types of extreme events (e.g., wildfires and earthquakes) could be achieved by adding relevant features and retraining the model on new data.

### v2

@@
 Incorporating social and demographic indicators reveals underlying patterns of community vulnerability and provides a clearer understanding of outage risk during extreme conditions.
-Four machine learning models—Random Forest (RF), Support Vector Machine (SVM), Adaptive Boosting (AdaBoost), and Long Short-Term Memory (LSTM) —are evaluated.
+Four machine learning models—Random Forest (RF), Graph Neural Network (GNN), Adaptive Boosting (AdaBoost), and Long Short-Term Memory (LSTM) —are evaluated.
 Experimental validation is performed on a large-scale dataset covering counties in the lower peninsula of Michigan.
-Among all models tested, the LSTM network achieves the lowest prediction error.
-Additionally, the result demonstrates that stronger economic conditions and more developed infrastructure are associated with lower outage occurrence.
+Among all models tested, the LSTM network achieves higher accuracy.
 keywords:
-Power system resilience, power outage prediction, machine learning.
+Power system resilience, power outage prediction, machine learning, extreme events.
 1 Introduction
-Low-Probability High-Consequence (LPHC) events—such as floods, hurricanes, and winter storms—pose severe threats to power systems reliability and resilience.
-[1].
-These events can cause large-scale blackouts and economic losses, as seen in Hurricane Sandy (2012), Irma (2017), and the 2021 Texas winter storm [2, 3, 4].
-Several studies has explored a wide range of modeling approaches to understand and predict the relationship between extreme weather conditions and outage occurrence.
+High-Impact Low-Probability (HILP) events—such as floods, hurricanes, and winter storms—pose severe threats to power systems reliability and resilience.
+xu2023review .
+These events can cause large-scale blackouts and economic losses, as seen in Hurricane Sandy (2012), Irma (2017), and the 2021 Texas winter storm buzzfeed2022 , texastribune2021 , daeli2022power .
+Several studies have explored a wide range of modeling approaches to understand and predict the relationship between extreme weather conditions and outage occurrence.
 For example, Eskandarpour et al.
-[5] predicted component damage prior to hurricanes by establishing decision boundaries using logistic regression.
-Spatial–temporal poisson regression has also been applied to model transmission line failure probabilities during severe weather events [6].
+R13 predicted component damage prior to hurricanes by establishing decision boundaries using logistic regression.
+Spatial–temporal poisson regression has also been applied to model transmission line failure probabilities during severe weather events yang2019 .
 Although these methods offer clear interpretability, their linear assumptions limit their ability to capture the complex multi-variable interactions.
-ML models provide an effective framework to manage complex non-linear relationships and process substantial data volumes in real-time [7].
+ML models provide an effective framework to manage complex non-linear relationships and process substantial data volumes in real-time N30 .
 These models outperform traditional statistical approaches due to their ability to model nonlinearities and integrate diverse datasets, including weather, infrastructure, and socio-economic indicators.
-Classification-based ML studies such as [8, 9] used models like SVM and boosted classifiers to determine outage occurrence or component damage.
-Regression-based studies have used tree-based ensembles such as Decision Trees, RF, AdaBoost, and XGBoost to predict outage counts, durations, or affected customers [10, 11, 12].
+Classification-based ML studies such as N3 , eskandarpour2017leveraging used models like SVM and boosted classifiers to determine outage occurrence or component damage.
+Regression-based studies have used tree-based ensembles such as Decision Trees, RF, AdaBoost, and XGBoost to predict outage counts, durations, or affected customers N1 , cruz , N41 .
 Deep learning approaches further extend predictive capability by capturing temporal and sequential patterns in outage and weather data.
 Arif et al.
-[13] employed deep neural networks to estimate restoration times, and recurrent architectures such as recurrent neural networks (RNN) to model temporal outage progression under varying environmental conditions.
+N4 employed deep neural networks to estimate restoration times, and recurrent architectures such as recurrent neural networks (RNN) to model temporal outage progression under varying environmental conditions.
 These models offer strong predictive performance but require large, high-quality datasets.
@@
 Jaech et al.
-[14] applied RNN-based models with environmental and operational data to predict outage duration in Seattle, while Sun et al.
-[15] used AdaBoost to estimate outage counts in Kansas.
+R14 applied RNN-based models with environmental and operational data to predict outage duration in Seattle, while Sun et al.
+R12 used AdaBoost to estimate outage counts in Kansas.
 Sarwat et al.
-[16] showed that day-to-day weather conditions influence power outage.
-Our work focuses on LPHC events that generate large-scale disruptions and require fundamentally different modeling considerations.
+sarwat2016 showed that day-to-day weather conditions influence power outage.
+Our work focuses on HILP events that generate large-scale disruptions and require fundamentally different modeling considerations.
 The state-of-the-art studies often encounter challenges that can compromise the accuracy of their analyses.
@@
 In particular, during such events, several monitoring devices may fail, resulting in incomplete and imbalanced datasets.
-In [17, 18, 19], Synthetic Minority Oversampling Technique (SMOTE) has been used to address data imbalance in classification tasks in power outage studies.
-However, SMOTE is effective for categorical targets, it is less productive for LPHC outage prediction, which is framed as a regression problem.
-Another limitations arises when models are trained on a single extreme event, as they exhibit limited adaptability to concurrent incidents with complex, multi-variable interactions [20].
-As shown in [21], combining data from two winter storm events instead of one improves power outage prediction.
+In N26 , N14 , N17 , Synthetic Minority Oversampling Technique (SMOTE) has been used to address data imbalance in classification tasks in power outage studies.
+However, SMOTE is effective for categorical targets, it is less productive for HILP outage prediction, which is framed as a regression problem.
+Another limitations arises when models are trained on a single extreme event, as they exhibit limited adaptability to concurrent incidents with complex, multi-variable interactions N15 .
+As shown in N16 , combining data from two winter storm events instead of one improves power outage prediction.
 Furthermore, socio-economic features are often missing in the previous models.
 As demonstrated in Cruz et al.
-[11], socio-economic indicators such as income levels and housing characteristics provide important context for understanding community vulnerability.
+cruz , socio-economic indicators such as income levels and housing characteristics provide important context for understanding community vulnerability.
 Counties with lower income or higher unemployment tend to have fewer resources for preparedness, which can increase outage severity and prolong restoration times.
-Similar observations have been reported in other studies [13, 14], where result shows that socio-economic factors contribute to prolonged outage durations and higher customer outage number.
-Incorporating socio-economic variables therefore enables a more comprehensive understanding of outage dynamics, supports improved prediction performance for LPHC events.
-In our prior work [19], we developed ML models to predict power outage probabilities, investigating the impact of socio-economic and infrastructure factors on non-LPHC outage events.
+Similar observations have been reported in other studies N4 , R14 , where result shows that socio-economic factors contribute to prolonged outage durations and higher customer outage number.
+Incorporating socio-economic variables therefore enables a more comprehensive understanding of outage dynamics, supports improved prediction performance for HILP events.
+In our prior work N17 , we developed ML models to predict power outage probabilities, investigating the impact of socio-economic and infrastructure factors on non-HILP outage events.
 This paper builds on those findings and proposes enhanced methodologies to predict power outages due to extreme weather events.
 The contributions of this paper are outlined as follows:
-1) Comprehensive Data Integration: Our model combines weather, socio-economic, and infrastructure data, along with seasonal indicators, to capture the complex factors influencing power outages.
+1) This model combines dynamic data such as weather, seasonal indicators, historical power outages with static data such as socio-economic, and infrastructure data to capture the complex factors influencing power outages.
 This comprehensive integration significantly enhances the predictive performance of the model.
-2) Innovative Use of SMOGN and KNN to Address Data Scarcity in LPHC Events: This work introduces the application of the Synthetic Minority Over-sampling Technique for Regression (SMOGN) and K-Nearest Neighbors (KNN) methods to mitigate data imbalance and missing data challenges associated with LPHC events.
-By adapting these techniques to the outage prediction context, the proposed framework effectively addresses data scarcity and improves the accuracy and robustness of LPHC event prediction.
-3) Evaluation of ML Models: We evaluate multiple ML and deep learning (DL) models, with LSTM proven particularly effective due to its ability to capture temporal patterns in time-series data.
+2) This work introduces the application of the Synthetic Minority Over-sampling Technique for Regression (SMOGN) and K-Nearest Neighbors (KNN) methods to mitigate data imbalance and missing data challenges associated with HILP events.
+By adapting these techniques to the outage prediction context, the proposed framework effectively addresses data scarcity and improves the accuracy and robustness of HILP event prediction.
+3) We evaluate and compare multiple ML and deep learning (DL) models for extreme event outage prediction using integrated static and dynamic data.
 This makes our model highly suited for real-world power outage forecasting.
-4) Large-Scale Validation: We validate our approach on a large-scale dataset spanning 2014 to 2022 across Michigan counties, ensuring the model’s robustness and generalizability to diverse geographic regions and socio-economic conditions.
+4) We train our approach on a large-scale power outage dataset across several Michigan counties.
+The dataset includes historical extreme events from 2014 to 2024, including the flooding event in Wayne county in June, 2020.
+This model can be generalized to diverse geographic regions and socio-economic conditions.
 The remainder of the paper is structured as follows: Section II presents the data description and preparation.
@@
 2.1 Historical Outage Data
-In this study, the EAGLE-I dataset is utilized to analyze severe power outages caused by LPHC events.
-The EAGLE-I dataset [22], collected and managed by Oak Ridge National Laboratory (ORNL), spans from November 2014 to December 2024.
+In this study, the EAGLE-I dataset is utilized to analyze severe power outages caused by extreme weather events.
+The EAGLE-I dataset N25 , collected and managed by Oak Ridge National Laboratory (ORNL) on behalf of DOE, spans from November 2014 to December 2024.
 This dataset, updated every 15 minutes, aggregates data from utility companies’ public outage maps using an ETL (Extract, Transform, Load) process to estimate the number of customers affected by power outages in each county.
@@
 1 illustrates the daily number of customer outages in Wayne County throughout 2024.
+In this context, a ’customer’ is defined as any entity that purchases energy from a utility under a specific tariff, rather than individual persons.
 Figure 1: Number of customers outage in Wayne County during 2024 based on EAGLE-1 dataset.
-In this context, a “customer” is defined as any entity that purchases energy from a utility under a specific tariff, rather than individual persons.
-Given that in the EAGLE-I, number of customers affected by power outage is recorded, and to ensure a fair comparison between counties with different sizes, it is crucial to adjust it for population differences.
-This can be done by collecting population data for each county from the U.S.
-Census Bureau [^2] https://data.census.gov/table.
-This approach allows us to normalize outage data relative to population size.
-In this way, the number of customers in the EAGLE-I dataset is divided by the population of the county scaled by 2 as shown in (1).
-$P_{o}=\frac{Number\>of\>customers\>lost\>power\>*2}{Population\>of\>the\>county}$ (1)
-It is estimated that, on average, each customer corresponds to approximately two residents [23].
 2.2 Historical Weather Data
@@
 We have chosen 8 features including air temperature (F), precipitation (inch), wind speed (km/h), wind gusts (km/h), shortwave radiation ($W/m^{2}$), relative humidity (%), cloud cover (%), and surface pressure (hPa).
-The Fig.
-2 shows the locations of weather stations across the Lower Peninsula of Michigan used in this study.
-The historical weather data often has missing data, particularly during LPHC events. To address this issue, we employ K-Nearest Neighbors (KNN) imputation method describe as follows:
-Algorithm 1 KNN Algorithm for Finding the Nearest Counties
+The historical weather data often has missing data, particularly during HILP events.
+To address this issue, KNN imputation method has been applied biswas5151316data .
+For a given county, the $k$ closest counties are determined by calculating the Euclidean distance between county.
+Then, impute the missing data by averaging the selected nearest neighbors, identified by the KNN algorithm as follows in Algorithm 1.
+Algorithm 1 Algorithm to Estimate Missing Data
 Input:
-$k\leftarrow 5$ (Number of nearest counties to be found)
-$S_{i}\leftarrow\text{Station}_{i}$ (Find the k nearest counties of $S_{i}$)
-$(x_{i},y_{i})$ (Coordinates of $S_{i}$)
-$D_{ij}$ (Euclidean distance between $S_{i}$ and $S_{j}$)
+Set of counties $\mathcal{C}=\{C_{1},C_{2},\dots,C_{N}\}$
+Number of nearest counties $k$ (set to $k=5$)
+Geographic coordinates $(x_{i},y_{i})$ of county $C_{i}$
+$D_{ij}$ (Euclidean distance between $C_{i}$ and $C_{j}$)
 Procedure:
-for $k=1,\dots,j$ do
-      Calculate $D_{ij}$ between $S_{i}$ and $S_{j}$ using equation (2)
-      Sort distances between $S_{i}$ and all other counties in descending order
-      Select the first $k$ closest counties to $S_{i}$
+for each county $C_{i}\in\mathcal{C}$ do
+  Initialize distance list $\mathcal{D}_{i}\leftarrow\emptyset$
+  for each county $C_{j}\in\mathcal{C}(i\not=j)$ do
+   Compute distance:
+$D_{ij}=\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}}$
+   Store $(C_{j},D_{ij})$ to $\mathcal{D}_{i}$
   end for
-Output: 5 nearest counties to $S_{i}$
-For a given $county_{i}$, the $k$ closest counties are determined by calculating the Euclidean distance between $county_{i}$ and every other county, as described in (2).
-$\text{Euclidean Distance}(S_{i},S_{j})$ $=\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}}$ (2)
-where $(x_{i},y_{i})$ and $(x_{j},y_{j})$ represent the coordinates of counties $S_{i}$ and $S_{j}$, respectively. Then, impute the missing data by averaging the selected neighbors:
-$Missing\>Value=\frac{1}{j}\sum_{i=1}^{j}{Observed\>value_{i}}$ (3)
-where $j$ represents the number of nearest counties considered for the imputation, identified by the K-Nearest Neighbors (KNN) algorithm [24]. Finally, the dataset is normalized using min-max scaler.
+  Sort $\mathcal{D}_{i}$ in ascending order of $D_{ij}$
+  Select the first $k$ counties from $\mathcal{D}_{i}$
+  Compute the average value of the $k$ neighbors
+end for
+Output: Impute the missing data
 Figure 2: Weather stations locations in Lower Peninsula of Michigan.
@@
 For this study, socio-economic indicators were obtained from the U.S.
-Census Bureau, including the distribution of the year when residential structures were built, the average household income, and the county-level unemployment rate.
+Census Bureau, including the distribution of the year when residential structures were built (2021 estimate), the average household income (2021 estimate), and the county-level unemployment rate (2021 estimate).
 The selected data was retrieved from the U.S.
-Census Bureau American Community Survey (ACS) 5-Year Estimates via the Census API.
+Census Bureau’s 2021 American Community Survey (ACS) 5-Year Estimates via the Census API.
 [^4] https://api.census.gov/data/2021/acs/acs5
@@
 2.4 Power Infrastructure Data
-The infrastructure features, including the number of poles, towers, substations, transformers, and lines for each county, are collected from the Open Street Map dataset [25] as shown in Fig.
+The infrastructure features, including the number of poles, towers, substations, transformers, and lines for each county, are collected from the Open Street Map dataset N19 as shown in Fig.
 3.
+Power infrastructure data provides valuable insights into the characteristics of each county.
 Given that the detailed topology of the power distribution network is typically not publicly available, data from the Open Street Map can serve as a useful approximation of the actual topology for research purposes.
-Furthermore, they provide information on how extensively the system might be affected by extreme conditions such as high winds [26].
+Furthermore, they provide information on how extensively the system might be affected by extreme conditions such as high winds N6 .
 For this study, the total number of each feature within a county is aggregated and then counted to provide a comprehensive overview of the area’s characteristics.
@@
 This method allows us to determine which county has a higher proportion of a particular infrastructure component, such as towers.
-Figure 3: Power infrastructure network of counties in Lower Peninsula of Michigan.
-3 Data-driven Outage Prediction Algorithm
-The proposed multi-step county-level power outage prediction algorithm is illustrated in Fig.
+Figure 3: Power infrastructure network of 20 counties in Lower Michigan.
+3 Methodology
+The proposed multi-step county-level power outage prediction algorithm due to extreme weather events is illustrated in Fig.
 4.
@@
 Second, the integrated dataset is rebalanced to reduce the impact of data imbalance and missing data.
-Third, multiple machine learning models are trained to capture the complex relationships between weather, socio-economic, and infrastructure factors and outage occurrence.
+Third, multiple machine learning models are trained on the data to capture the complex relationships between weather, socio-economic, and infrastructure factors and outage occurrence.
 Finally, the performance of the models is evaluated using standard metrics.
-Figure 4: The framework of the proposed power outage prediction algorithm.
+In the following, we discuss the last three steps, and then the model formulation is presented.
+Figure 4: The proposed learning-based framework for power outage prediction.
 3.1 Identifying Extreme Weather Event Data
@@
 Fig.
-5 illustrates the distribution of extreme weather events across 20 Michigan counties.
+5 illustrates the distribution of extreme weather events across 20 Michigan counties, showing that Wayne and Hillsdale counties experienced the highest number of events.
 Figure 5: Number of Extreme Events by County (2014–2024).
@@
 To address this limitation, seasonal information aligned with the timing of distinct extreme weather patterns is incorporated.
-For a specified extreme event day (denoted as $d_{ex}$), we identify all days within the same season and include them in the set $R$.
-This is defined as:
-$R=\{d_{1},d_{2},\ldots,d_{j}\}\text{ where }d_{j}\in(m_{i-1},m_{i},m_{i+1})$
-where, $m_{i}$ represents a month within a season, selected from the set $M=\{m_{1},m_{2},m_{3},\ldots,m_{12}\}$.
-The next step involves calculating the $L2norm$ of the weather features between the extreme day ($d_{ex}$) and all other days in set $R$.
-This comparison helps identify days with weather patterns most similar to the extreme event day.
-Assume $w_{d_{ex}}$ is the matrix of weather features for the extreme day (comprising 8 features, one for each weather features).
-The weather features for the $j$ days are represented by the matrix $W_{d}$ as follows:
-$W_{d_{ex}}=\begin{pmatrix}w_{1d_{ex}}&w_{2d_{ex}}&\cdots&w_{8d_{ex}}\\ \end{pmatrix}$
-$W_{d}=\begin{pmatrix}w_{1d_{1}}&w_{2d_{1}}&\cdots&w_{8d_{1}}\\ w_{1d_{2}}&w_{2d_{2}}&\cdots&w_{8d_{2}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{1d_{j}}&w_{2d_{j}}&\cdots&w_{8d_{j}}\\ \end{pmatrix}$ (4)
-The L2normL2\>norm between the row vector $w_{d_{ex}}$ of $W_{d_{ex}}$ and each row vector $w_{d_{j}}$ of $W_{d}$ is calculated using the following formula:
-$L2\>norm(w_{d_{ex}},w_{d_{j}})=\sqrt{\sum_{k=1}^{8}(w_{kd_{ex}}-w_{{k}d_{j}})^{2}}$ (5)
-Here, $W_{d_{ex}}=\begin{pmatrix}w_{1d_{ex}}&w_{2d_{ex}}&\cdots&w_{8d_{ex}}\\ \end{pmatrix}$ is the row vector of extreme day’s weather features, $W_{d_{j}}=\begin{pmatrix}w_{1d_{j}}&w_{2d_{j}}&\cdots&w_{8d_{j}}\\ \end{pmatrix}$ is the j-th row vector of $W_{d}$, $w_{kd_{ex}}$ and $w_{{k}d_{j}}$ are the elements of $W_{d_{ex}}$ and $W_{d_{j}}$, respectively.
-Subsequently, the calculation is repeated for each row in the matrix $W_{d}$, corresponding to each day $d_{j}$ within the set $R$.
-For each day, compute the L2normL2\>norm between the weather feature vector of the extreme day $W_{d_{ex}}$ and that of the day $d_{j}$, denoted as $W_{d_{j}}$.
-This results in a $j\times 1$ matrix where each element represents the $L2norm$ between $W_{d_{ex}}$ and $W_{d_{j}}$.
-The process is outlined mathematically as follows:
-$L2\text{Norm}=\begin{pmatrix}L2\>norm(W_{d_{ex}},W_{d_{1}})\\ L2\>norm(W_{d_{ex}},W_{d_{2}})\\ \vdots\\ L2\>norm(W_{d_{ex}},W_{d_{j}})\\ \end{pmatrix}$ (6)
-Finally, we select the weather features and outage data of days with the lowest $L2norm$ from matrix $L2Norm$, indicating the highest similarity to the extreme event day.
-Incorporating these seasonal references significantly enhances the model’s ability to differentiate between various outages in the dataset.
-Then, the dataset is rebalanced and used to train the machine learning models described in the following subsections.
-Algorithm 2 SMOGN for Outage-Ratio Rebalancing
-Input:
-    $D\leftarrow$ Training dataset (hourly), with features X
-    $T\leftarrow$ Threshold of minority and majority instances
-    $k\leftarrow$ Number of nearest neighbors
-    $O\%\leftarrow$ Percentage of over-sampling
-    $U\%\leftarrow$ Percentage of under-sampling
-Partition:
-    $S_{min}\leftarrow$ $y_{i}\geq T$ (Minority instances)
-    $S_{maj}\leftarrow$ $y_{i}<T$ (Majority instances)
-/*Over-Sampling minority instances*/
-for $S\in S_{min}$ do
-  $\hat{S}_{min}\leftarrow|S|\times O\%$
-  for $sample\in S$ do
-   $KList\leftarrow$ $k$ nearest neighbors of sample
-   $Distances\leftarrow$ distance between sample and each neighbor
-   $SafeRange\leftarrow{0.5\times\text{median}(Distances)}$
-   for $i=1$ to $\hat{S}_{min}$ do
-     $X\leftarrow$ randomly selected neighbor from KListKList
-     if $Distance(X,sample)<SafeRange$ then
-      $X_{new}\leftarrow$ SMOTER interpolation
-     else
-      $X_{new}\leftarrow$ samplesample + Gaussian Noise
-     end if
-     Append $X_{new}$ to $D_{new}$
-   end for
-  end for
-end for
-/*Under-Sampling majority cases*/
-for $S\in S_{maj}$ do
-  $under\_sampled\_instances\leftarrow$ randomly sample $|S|\times U\%$ instances
-  Append under_sampled_instancesunder\_sampled\_instances to $D_{new}$
-end for
-Output: Rebalanced training set $D_{new}$
+Since HILP events are characterized by large outage magnitudes, customer outage counts are used directly as the severity indicator.
+Two complementary indicators are defined to identify HILP outage conditions.
+Let $y(c,t)$ denote the number of customers without power in county $c$ at time $t$, as reported by the EAGLE-I dataset.
+Since HILP events are characterized by large outage magnitudes, customer outage counts are used directly as the severity indicator.
+3.1.1 Identification of HILP Outage Events
+Two complementary indicators are defined to identify HILP outage conditions.
+First, a weather-based extreme indicator is defined as
+[…]
+Second, an extreme outage indicator is defined using the empirical distribution of customer outage counts during extreme weather periods.
+Let $Q_{\alpha}$ denote the $\alpha$-quantile of $y(c,t)$.
+An outage is considered high-consequence if
+$I^{\mathrm{HC}}_{c,t}=\begin{cases}1,&y(c,t)\geq Q_{\alpha},\\ 0,&\text{otherwise},\end{cases}$ (2)
+where $\alpha=0.7$ is used in this study, corresponding to the top 30% of outage magnitudes observed during extreme weather conditions.
+The set of HILP seed events is therefore defined as
+$\mathcal{E}_{0}=\left\{(c,t)\mid I^{\mathrm{WX}}_{c,t}\cdot I^{\mathrm{HC}}_{c,t}=1\right\}.$ (3)
+These seed events represent extreme outage and serve as reference points for identifying meteorologically similar extreme conditions.
+3.1.2 Seasonal Analog Selection Using Weather Similarity
+For each HILP seed event $(c,t_{\mathrm{ex}})\in\mathcal{E}_{0}$, a set of candidate analog periods is selected to preserve seasonal consistency.
+Let $m(t)$ denote the calendar month associated with time $t$.
+The seasonal candidate set is defined as
+$\mathcal{R}(c,t_{\mathrm{ex}})=\left\{(c,t)\;\middle|\;m(t)\in\left\{m(t_{\mathrm{ex}})-1,\;m(t_{\mathrm{ex}}),\;m(t_{\mathrm{ex}})+1\right\}\right\}.$ (4)
+For each county and time, an aggregated weather feature vector is constructed as
+$\mathbf{x}_{c,t}=\big[T_{c,t},\,\mathrm{Prec}_{c,t},\,WS_{c,t},\,WG_{c,t},\,SWR_{c,t},\,RH_{c,t},\,CC_{c,t},\,SP_{c,t}\big]^{\top},$ (5)
+where the variables represent daily or hourly aggregates of temperature, precipitation, wind speed, wind gust, shortwave radiation, relative humidity, cloud cover, and surface pressure.
+To ensure balanced comparison across heterogeneous weather variables, feature-wise standardization is applied:
+$\tilde{\mathbf{x}}_{c,t}=\frac{\mathbf{x}_{c,t}-{\mu}}{{\sigma}},$ (6)
+where $\mu$} and $\sigma$} denote the mean and standard deviation vectors computed over the full weather dataset.
+Weather similarity between a seed event $t_{\mathrm{ex}}$ and a candidate time $t$ is quantified using the Euclidean distance in standardized feature space:
+$\delta(c,t_{\mathrm{ex}},t)=\left\|\tilde{\mathbf{x}}_{c,t_{\mathrm{ex}}}-\tilde{\mathbf{x}}_{c,t}\right\|_{2}.$ (7)
+For each seed event, the $K$ most similar analog periods are selected as
+$\mathcal{A}(c,t_{\mathrm{ex}})=\operatorname*{arg\,min}_{\begin{subarray}{c}t\in\mathcal{R}(c,t_{\mathrm{ex}})\\ |\mathcal{A}|=K\end{subarray}}\delta(c,t_{\mathrm{ex}},t).$ (8)
+The final extreme-event dataset used for training is constructed as
+$\mathcal{E}=\mathcal{E}_{0}\;\cup\;\bigcup_{(c,t_{\mathrm{ex}})\in\mathcal{E}_{0}}\mathcal{A}(c,t_{\mathrm{ex}}).$ (9)
+This procedure expands the HILP conditions by incorporating meteorologically similar extreme-weather periods while preserving seasonal characteristics.
+Algorithm 2 Data Rebalancing for High-Impact Customer Outages
+1:Input:
+2:    $\mathcal{T}\leftarrow$ Hourly training set with feature matrix $\mathbf{Z}$ and target $q$
+3:    $\tau\leftarrow$ Outage-impact threshold
+4:    $K\leftarrow$ Number of nearest neighbors
+5:    $\rho_{o}\leftarrow$ Over-sampling rate
+6:    $\rho_{u}\leftarrow$ Under-sampling rate
+7:Partition:
+8:    $\mathcal{C}_{\text{hi}}\leftarrow\{(\mathbf{z}_{j},q_{j})\in\mathcal{T}\mid q_{j}\geq\tau\}$ (High-impact cases)
+9:    $\mathcal{C}_{\text{lo}}\leftarrow\{(\mathbf{z}_{j},q_{j})\in\mathcal{T}\mid q_{j}<\tau\}$ (Low-impact cases)
+10:    $\mathcal{T}_{\text{bal}}\leftarrow\emptyset$
+11:/* Synthetic generation for high-impact outages */
+12:for $(\mathbf{z}_{r},q_{r})\in\mathcal{C}_{\text{hi}}$ do
+13:  $N_{\text{syn}}\leftarrow\rho_{o}$
+14:  $\mathcal{N}_{r}\leftarrow$ $K$ nearest neighbors of $(\mathbf{z}_{r},q_{r})$ in $\mathcal{C}_{\text{hi}}$
+15:  $\Delta_{r}\leftarrow$ pairwise distances between $(\mathbf{z}_{r},q_{r})$ and $\mathcal{N}_{r}$
+16:  $\lambda_{r}\leftarrow 0.5\times\text{median}(\Delta_{r})$
+17:  for $m=1$ to $N_{\text{syn}}$ do
+18:   $(\mathbf{z}_{m},q_{m})\leftarrow$ randomly selected neighbor from $\mathcal{N}_{r}$
+19:   if $\mathrm{dist}\big((\mathbf{z}_{m},q_{m}),(\mathbf{z}_{r},q_{r})\big)<\lambda_{r}$ then
+20:     $(\mathbf{z}^{*},q^{*})\leftarrow$ SMOTER interpolation between $(\mathbf{z}_{r},q_{r})$ and $(\mathbf{z}_{m},q_{m})$
+21:   else
+22:     $\mathbf{z}^{*}\leftarrow\mathbf{z}_{r}+\boldsymbol{\eta}$ (Gaussian perturbation)
+23:     $q^{*}\leftarrow q_{r}$
+24:   end if
+25:   Append $(\mathbf{z}^{*},q^{*})$ to $\mathcal{T}_{\text{bal}}$
+26:  end for
+27:end for
+28:/* Down-sampling low-impact outages */
+[…]
+30:Store $\mathcal{C}_{\text{lo}}^{\downarrow}$ to $\mathcal{T}_{\text{bal}}$
+31:Output: Balanced training dataset $\mathcal{T}_{\text{bal}}$
 3.2 Input Data Balancing
-A key challenge in predicting LPHC outages is the imbalance in outage data, where small-scale events dominate and large-scale extreme events are minority.
-This imbalance causes ML models to favor frequent low-impact cases, leading to poor performance on the critical but infrequent LPHC events.
-In order to balance the dataset, we apply SMOGN, which has been introduced to address the issue of imbalanced datasets in regression tasks [27].
-It integrates two over-sampling technique, SMOTER and SMOTER-GN [28].
-While the $L2norm$ filtering step extracts outages associated with extreme events, the SMOGN algorithm addresses the imbalance in outage severity within these events.
+A key challenge in predicting HILP outages is the imbalance in outage data, where small-scale events dominate and large-scale extreme events are minority.
+This imbalance causes ML models to favor frequent low-impact cases, leading to poor performance on the critical but infrequent HILP events.
+In order to balance the dataset, we apply SMOGN, which has been introduced to address the issue of imbalanced datasets in regression tasks N20 .
+It integrates two over-sampling technique, SMOTER and SMOTER-GN Nina25 .
+This algorithm addresses the imbalance in outage severity within these events.
 SMOGN generates synthetic data by applying three key techniques: random under-sampling, SMOTER (Synthetic Minority Over-sampling Technique for Regression), and adding Gaussian noise.
@@
 It generates new synthetic samples by interpolating rare cases.
-This is done by taking a weighted average of the target variable values between the original minority instance and its k-nearest neighbors [28].
+This is done by taking a weighted average of the target variable values between the original minority instance and its k-nearest neighbors Nina25 .
 Figure 6: Probability distribution of customer outages in the EAGLE-I dataset (2014–2024), shown before and after applying SMOGN.
 SMOGN is summarized in Algorithm 2: We mark a case as rare if its outage value is in the top 30% of the target distribution.
-A rarity threshold of 380 customers out was used and five nearest rare dataset neighbors were considered for SMOTER interpolation within a safe region.
+A rarity threshold of 380 customers out was used.
+five nearest rare dataset neighbors were considered for SMOTER interpolation within a safe region.
 Also, Gaussian noise with perturbation 2% was added when neighbors lie outside the local safe range.
@@
 6 shows the influence of SMOGN on the outage occurrence probability distribution for the extreme weather dataset used in the case study.
-The original dataset is dominated by small outages near zero, while the rebalanced dataset includes more high-consequence outages.
+The original dataset is dominated by small outages near zero, while the rebalanced dataset includes more high-consequence outages, leading to a smoother and more balanced distribution for training.
 After rebalancing, probability is reduced in the low-outage region and increased in the mid/high-outage region beyond the threshold.
@@
 The model is trained with 100 decision trees and evaluated on the same dataset.
-Support Vector Machine (SVM): SVM for regression identifies an optimal hyperplane in a higher-dimensional space that best fits the data while maximizing the distance from the nearest points.
-It minimizes errors, making it suitable for predicting continuous variables.
-SVM uses kernel functions to map the data into a space where these nonlinear patterns become easier to fit.
+Graph Neural Network (GNN): GNNs are deep learning architectures designed to operate on graph-structured data LiaoMPCE .
+In this study, counties are represented as nodes in a spatio-temporal graph, with edges connecting geographically proximate counties (spatial edges) and linking the same county across consecutive time steps (temporal edges).
+Each node is characterized by weather, socio-economic, and infrastructure features.
+A Graph Attention Network (GAT) architecture is employed to perform message passing, where the attention mechanism adaptively weights the influence of neighboring nodes, enabling the model to focus on the most relevant spatial-temporal relationships.
+The model follows the procedure outlined in Algorithm 3.
+Algorithm 3 Spatio-Temporal Outage Prediction Using Graph Attention Network (GAT)
+1:Input: $D\leftarrow$ Dataset with hourly weather, socio-economic, spatial and extreme events data
+2:Target: $y\leftarrow$ Number of customers outage during extreme events
+3:Step 1: Preprocessing
+4:Normalize features $\mathcal{F}$, target $y$ using Min-Max scaling
+5:Step 2: Node define
+6:Assign node IDs based on (county, time)
+7:Step 3: Graph Construction
+8:Spatial edges: For each timestamp $t$, connect counties within 50 miles
+9:Temporal edges: For each county $c$, link consecutive timestamps $t_{i}\rightarrow t_{i+1}$
+10:Step 5: Two Layers GAT Model
+11:Layer 1: 6464-dim GAT with 44 heads + ReLU
+12:Layer 2: 3232-dim GAT + ReLU
+13:Output: Linear layer for scalar prediction
+14:Step 6: Model Training
+15:for epoch =1=1 to 150150 do
+16:  Forward pass: $\hat{y}=\texttt{GAT}(G)$
+17:end for
+18:Step 7: Post-processing
+19:Predict scaled output $\hat{y}_{\text{scaled}}$
+20:Inverse-transform to obtain actual outage: ${y}=\texttt{InverseScaler}(\hat{y}_{\text{scaled}})$
+21:Output: Predicted customer outages $y$
 Adaptive Boosting (AdaBoost): It is a boosting algorithm for regression which strengthens weak models by prioritizing harder-to-predict instances, refining the ensemble’s accuracy with each iteration.
 The final prediction is a weighted sum of the predictions from all weak learners, with weights determined based on their performance.
-Given the imbalanced nature of outage data, AdaBoost can concentrate learning on rare and high-impact outage conditions.
 Long Short-Term Memory (LSTM): LSTM networks are a type of RNN specialized in processing sequences of data.
@@
 3.4 Performance Evaluation
-The following common performance metrics are used for models evaluation [29]:
-Mean Absolute Error (MAE): It measures the average deviation of the predicted values from the actual observed values. It shows how far the model’s predictions are from the actual values on average.
-[…]
-Mean Square Error (MSE): That metric measures the average of the squared differences between the estimated values and the actual observed values.
-Lower MSE indicates that the model is more consistent and accurate across the full range of observations.
-$MSE=\frac{1}{n}\sum_{i=1}^{n}(P_{o}(i)-\hat{P_{o}(i)})^{2}$ (8)
-In both expressions, $P_{o}(i)$ denotes the actual observed value and $\hat{P_{o}(i)}$ represents the predicted value.
+The following common performance metrics for regression tasksplevris2022investigation are used for models evaluation:
+Mean Absolute Percentage Error (MAPE): It quantifies the average absolute deviation between predictions and observations as a percentage of the observation.
+$\mathrm{MAPE}(\%)=\left(\frac{1}{n}\sum_{i=1}^{n}\left|\frac{P(i)-\hat{P}(i)}{P(i)}\right|\quad\right)\times 100$ (10)
+Lower values indicate better performance. For example, a MAPE of 8% implies an average error equal to 8%.
+$R^{2}$ Score (Coefficient of Determination): It quantifies the proportion of variance in the observed data that is explained by the model’s predictions.
+An $R^{2}$ value of 100% indicates perfect prediction, while a value of 0 implies that the model performs no better than predicting the mean of the observed values.
+$R^{2}=\left(1-\frac{\sum_{i=1}^{n}\left(P(i)-\hat{P}(i)\right)^{2}}{\sum_{i=1}^{n}\left(P(i)-\bar{P}\right)^{2}}\right)\times 100$ (11)
+where $P(i)$ is the actual outage value at instance $i$, $\hat{P}(i)$ is the predicted value, $\bar{P}$ is the mean of the actual values, and $n$ is the number of instances.
 3.5 Outage Prediction Model Formulation
-With all relevant features integrated into the dataset, the problem is formulated as predicting county-level power outages under extreme events. Eq. 3.5 shows the model formulation.
-$P_{o}(t)$ $=f\big(P_{o}(t-1),P_{o}(t-2),\cdots,P_{o}(t-n),$
-$\quad W(t),W(t-1),\cdots,W(t-n),S,I_{f}\big)$ (9)
-Here, $P_{o}(t)$ denotes the ratio of customers without power in a county at time $t$ during an extreme event, and $P_{o}(t-k)$ is the corresponding ratio observed during the $k$-th most recent extreme event.
-$W(t)$ is the vector of contemporaneous weather features for the county, while W(tk)W(t-k) contains the weather features recorded for the $k$-th most recent event.
-$S$ comprises socio-economic indicators, including average household income, unemployment rate, and the distribution of year built for residential structures.
-$I_{f}$ aggregates power-system attributes such as counts of poles, towers, substations, transformers, and line segments by county.
+With all relevant features integrated into the dataset, the problem is formulated as predicting county-level power outages under extreme weather events.
+Let $y(c,t)$ denote the number of customers without power in county $c$ at time $t$ during an identified extreme-weather period.
+Eq.
+12 shows the model formulation.
+$y(c,t)=f_{\theta}\!\left(\mathbf{Y}_{c,t}^{(n)},\;\mathbf{W}_{c,t}^{(n)},\;\mathbf{S}_{c},\;\mathbf{I}_{c}\right),$ (12)
+where $f_{\theta}(\cdot)$ denotes a data-driven mapping with parameters $\theta$ learned from historical data.
+The temporal outage history is represented as
+$\mathbf{Y}_{c,t}^{(n)}=\big[y(c,t-1),\;y(c,t-2),\;\ldots,\;y(c,t-n)\big],$
+and the corresponding historical weather features are given by
+$\mathbf{W}_{c,t}^{(n)}=\big[\mathbf{W}_{c,t},\;\mathbf{W}_{c,t-1},\;\ldots,\;\mathbf{W}_{c,t-n}\big],$
+where $\mathbf{W}_{c,t}$ denotes the vector of contemporaneous weather variables.
+The vector $\mathbf{S}_{c}$ contains county-level socio-economic indicators, including average household income, unemployment rate, and residential building age distribution, while $\mathbf{I}_{c}$ represents power infrastructure attributes.
+This formulation enables the predictive models to capture nonlinear temporal dependencies in outage evolution, the influence of extreme weather dynamics, and the modulating effects of infrastructure and socio-economic vulnerability on outage severity.
 4 Experimental Setup and Simulation Results
-The dataset is divided into training (2014 to 2018), validation (2019), and test (2020–2021) sets to evaluate the performance of the models.
-The implementation of RF, SVM, and AdaBoost is done using Python Scikit-learn library [30] and LSTM is implemented using Tensorflow library [31].
-We conducted hyperparameter tuning utilizing the Grid- Search tool in Python [30], employing cross-validation to identify the optimal input parameters for the models, as shown in Table 1.
-The LSTM network is a one-layer LSTM with 128 nodes followed by a Dense layer.
-The LSTM layer employs its standard activation structure, where the cell state is updated using a tanh activation and the gating mechanisms use sigmoid functions.
-Adam optimizer is used with a learning rate of 0.001 and a batch size of 128.
-Table 1: Cross-Validation Hyper-Parameters Selection
-Hyper-Parameter RF SVM AdaBoost
-Kernel rbf
-Learning Rate - - 0.001
-Estimators 4 - 120
-Loss square
-C - 10 -
-Max Depth 100 - -
-Gamma - 0.001 -
-Fig.
-7 illustrates the correlation analysis between socio-economic, infrastructure features and power outages.
-Higher mean income, lower unemployment rate, and more extensive infrastructure, including towers, lines, transformers, and poles, are associated with lower outages, suggesting that wealthier counties or those with more infrastructure might have higher power systems resilience.
-Substations and the year of structure construction also show a slight correlation, indicating newer or more substations may lead to improved outage resistance.
-Figure 7: Pearson’s correlation of the customers outage number with the infrastructure and socio-economic variables.
-We also conducted a feature importance analysis using the RF-based Mean Decrease in Impurity (MDI).
+The dataset spanning from 2014 to 2024 is used to train the model, while the June 6–8, 2020 flood event in Wayne County is selected for testing the performance.
+The implementation of RF and AdaBoost is done using Python Scikit-learn library N22 and LSTM is implemented using Tensorflow library N21 .
+The GNN model employed in this study is based on the GAT architecture using PyTorch Geometric.
+Hyperparameter tuning was performed using the GridSearch tool in Python N22 , with cross-validation employed to determine the optimal input parameters for each model, as summarized in Table 1.
+Table 1: Cross-Validation Hyper-Parameters Selection.
+Hyper-Parameter RF AdaBoost GNN LSTM
+Learning Rate - 0.001 0.01 0.001
+Estimators 100 120 - -
+Hidden Units - - 128 $\rightarrow$ 32 128
+Activation - - ReLU sigmoid
+Optimizer - - Adam Adam
+Max Depth - - - -
+Epochs - - 300 100
 Fig.
 7 depicts the contribution of each feature to the model’s predictive accuracy.
-The analysis showed that weather variables dominate outage prediction, with precipitation (19%), wind speed (13.5%), and surface pressure (13.1%) being the top contributors.
+The analysis result showed that weather features, such as precipitation and average wind speed, have the most significant contribution to the model’s performance.
+The analysis confirms that weather variables dominate outage prediction, with precipitation (19%), wind speed (13.5%), and surface pressure (13.1%) being the top contributors.
 Non-weather features, including infrastructure density (3.5%) and socioeconomic indicators such as average income (1.4%), also exhibit contribution to the prediction.
-Figure 8: Features importance across weather, infrastructure, and socioeconomic variables.
-After these relationships are established, the models are trained using different combinations of input data, both with and without the inclusion of socioeconomic, infrastructure, and weather features, so that the influence of each data layer on predictive performance can be examined.
-By comparing model accuracy across these configurations, the added value of these additional features is evaluated, and their overall impact on the outage prediction process is determined.
-Scenario-1 $=\big(P_{o}(t-1),P_{o}(t-2),\cdots,P_{o}(t-n),$
-$\quad W(t),W(t-1),\cdots,W(t-n)\big)$
-Scenario-2 $=\big(P_{o}(t-1),P_{o}(t-2),\cdots,P_{o}(t-n),$
-$\quad W(t),W(t-1),\cdots,W(t-n),S,I_{f}\big)$
-Scenario 1 uses weather data from the event day and the preceding day, together with outage information from the day before the extreme event and seasonal weather patterns.
-This configuration enables the models to utilize historical outage behavior along with weather conditions to estimate future outages.
-Scenario 2 expands upon this by incorporating all data used in Scenario 1 together with infrastructure and socioeconomic information.
-Table 2: Performance Comparison of Models in Scenario-1 and Scenario-2.
-Model Scenario-1 Scenario-2
-MSE MAE MSE MAE
-RF 0.0231 0.0898 0.0222 0.0827
-SVM 0.0175 0.0796 0.0172 0.0774
-AdaBoost 0.0210 0.0735 0.0204 0.0730
-LSTM 0.0127 0.0749 0.0086 0.0432
-Figure 9: Performance comparison between models in terms of MSE.
-Table.
-2 presents the performance of each model and shows that the scenario incorporating socioeconomic and infrastructure information produces lower prediction error.
-This outcome provides evidence that incorporating socioeconomic and infrastructure variables enhances the model’s predictive capability.
+Figure 7: RF features importance across weather, infrastructure, and socioeconomic variables.
+(a) Random Forest prediction vs. actual outages
+(b) GNN prediction vs. actual outages
+(c) AdaBoost prediction vs. actual outages
+(d) LSTM prediction vs. actual outages
+Figure 8: Comparison of predicted and actual outage counts for flood event in Wayne County (June 6–8, 2020).
 Fig.
-10 illustrates the performance of the model by comparing the predicted outage values with the actual recorded values for LSTM model.
-Each point represents a data sample from the validation period, and the diagonal line indicates the ideal one to one correspondence between predicted and observed outages.
-As shown in the figure, the majority of the predictions follow the general upward trend of the reference line.
-Figure 10: Scatter-plots of predicted versus actual customer outage for LSTM model.
-Fig.
-11 illustrates the comparison between the predicted and actual customer outage counts for the June 6–8, 2020 flood event in Wayne County, Michigan using four models: RF, SVM, AdaBoost, and LSTM.
+8 illustrates the comparison between the predicted and actual customer outage counts for the June 6–8, 2020 flood event in Wayne County, Michigan using four models: RF, GNN, AdaBoost, and LSTM.
 The time series plots show how well each model follows the temporal evolution and magnitude of the observed outages.
@@
 The Random Forest model also exhibits good performance.
-The AdaBoost model shows moderate agreement with the actual outages.
-The SVM model presents the weakest performance among the four.
+The GNN model shows moderate agreement with the actual outages.
+The AdaBoost model presents the weakest performance among the four.
+Fig.
+9 displays the performance of each model.
+Fig.
+9 confirming that LSTM (highest $R^{2}$) offers the most accurate predictions for outage events in this scenario.
+The RF and GNN model presents a promising deep learning-based approach with potential for further enhancement, especially for spatial-temporal generalization.
+Figure 9: Performance comparison between models in terms of Error (MAPE) and Accuracy ($R^{2}$).
+Table 2: Comparison of Proposed Work with State-of-the-Art Studies.
+Study Dataset Features Used Models Applied Prediction Setup and Results
+Lee et al.
+N41 EAGLE-I outage data Historical weather RF, KNN, XGBoost.
+XGBoost achieved the best performance.
+The study forecasts state-level maximum outages within 12 hours of a weather alert; however, it does not offer high-resolution predictions of outages caused by extreme events.
+Cruz et al.
+cruz EAGLE-I + NASA (weather), Census Bureau (socio-economic), FEMA Risk Index Weather + Socio-economic Regression, RF, XGBoost, KNN.
+RF outperform other models.
+The model predicts next-day average outages during Hurricane Matthew (Oct 2016).
+It does not offer high-resolution predictions.
+Our Work (Proposed) EAGLE-I + Open-Meteo (weather)+ Census (socio-economic), Infrastructure data Weather + Socio-economic + Infrastructure + Seasonal indicators RF, AdaBoost, LSTM, GNN.
+LSTM is the best performer.
+The model achieves county-level validation with hourly resolution in predicting affected customers during the June 2020 Wayne County flood.
 The performance of our proposed outage prediction framework is also compared with state-of-the-art studies in Table 2.
@@
 In contrast, our method integrates weather, socio-economic, infrastructure, and seasonal indicators to deliver accurate county-level, hourly outage predictions during extreme events.
-(a) RF prediction vs. actual outages
-(b) SVM prediction vs. actual outages
-(c) AdaBoost prediction vs. actual outages
-(d) LSTM prediction vs. actual outages
-Figure 11: Comparison of predicted and actual outage counts for the June 2020 flood event in Wayne County.
-Table 3: Comparison of Proposed Work With State-of-the-Art Studies.
-Study Dataset Features Used Best Model Prediction Resolution
-Lee et al. [12] EAGLE-I + NWS alerts Weather features only XGBoost
-($R^{2}$ = 0.998, RMSE = 2115 customers)
-Event-level maximum outage prediction;
-No hourly or county-level resolution
-Cruz et al. [11] EAGLE-I + NASA weather + Census data Weather + Socio-economic RF: MSE = 5.93e-05
-MAE = 0.00102
-RMSE = 0.00770
-Next-day average outage prediction;
-Not high-resolution
-Our Work (Proposed) EAGLE-I + Open-Meteo + Census + Infrastructure Weather + Socio-economic + Infrastructure + Seasonal indicators LSTM
-(MSE = 0.0086, MAE = 0.0432)
-County-level, hourly outage prediction during extreme events
 Given that this model is primarily trained on data from extreme weather and socio-economic features, it is well-suited to be used for planning to enhance the resilience of power systems.
 Planning can be done through simulating the system’s performance across a range of severe weather scenarios.
-5 Conclusions
-In this paper, we presented a novel machine learning–based approach for predicting power outages caused by LPHC events.
-The method integrated historical outage data from the EAGLE-I dataset, weather data from Open-Meteo, as well as infrastructure and socio-economic information to construct predictive models.
-We addressed the challenges of limited data availability and missing records for LPHC events by applying KNN and SMOGN methods.
-We showed the correlations between socio-economic and infrastructure features with extreme power outages.
-Several machine learning models, including RF, SVM, AdaBoost, and LSTM networks, were compared.
-The experimental evaluation on Michigan counties demonstrated that the LSTM model consistently outperformed RF, AdaBoost, and SVM in terms of lowest error.
-Our analysis indicated that weather-related variables—particularly precipitation and wind speed were the primary drivers of outages, while infrastructure and socio-economic factors exerted secondary influences.
-This model can be adapted to other regions by incorporating localized data, including region-specific weather patterns and infrastructure conditions as well as local social-economic information.
-Similarly, extending the model to handle other types of extreme events (e.g., wildfires and earthquakes) could be achieved by adding relevant features and retraining the model on new data.
+By running these simulations, planners can assess vulnerabilities, analyze system responses, and make data-driven decisions on where to allocate resources effectively to enhance grid resilience under severe conditions.
+5 Conclusion
+In this paper, we presented a novel machine learning–based approach for predicting power outages caused by HILP events.
+The proposed method integrated historical outage data from the EAGLE-I dataset, weather data from Open-Meteo, as well as infrastructure and socio-economic information to construct predictive models.
+We addressed the challenges of limited data availability and missing records for HILP events by applying KNN and SMOGN methods.
+Several machine learning models, including RF, GNN, AdaBoost, and LSTM networks, were compared.
+These models were trained and tested across scenarios with varying complexity.
+The experimental evaluation on 20 Michigan counties demonstrated that the LSTM model consistently outperformed RF, GNN and AdaBoost.
+Our analysis indicates that weather-related variables—particularly precipitation, wind speed, and surface pressure—are the primary drivers of outages, while infrastructure and socio-economic factors exert secondary influences.
+The model can be adapted to other regions by incorporating localized data, including region-specific weather patterns and infrastructure conditions and local social-economic information.
+Similarly, extending the model to handle other types of extreme events (e.g., wildfires and storms) could be achieved by adding relevant features and retraining the model on new data.
+Acknowledgment
+This paper is supported by the Department of Energy, Solar Energy Technologies Office (SETO) Renewables Advancing Community Energy Resilience (RACER) program under Award Number DE-EE0010413.
+The authors acknowledge early-stage discussions related to the doctoral dissertation research of Nina Fatehi.
+The content of this material represents the views of the authors only and is not endorsed by or necessarily reflective of the views of the Department of Energy.